OpenStudy (anonymous):

HELP!!! URGENT!!! For -1 ≤ x ≤, F(x) = integral of the square root of 1 - t^2 dt from -1 to x. a. What does F(1) represent geometrically? b. Find F'(x).

OpenStudy (anonymous):

\[\int\limits_{-1}^{x}\sqrt{1-t^{2}}dt\]

OpenStudy (anonymous):

 \[F(x)=\int_{-1}^x\sqrt{1-t^2}~dt\] So, \[F(1)=\int_{-1}^1\sqrt{1-t^2}~dt\] This definite integral gives you the area under the curve \(f(t)=\sqrt{1-t^2}\). What does that give you?